# SpecReason: Fast and Accurate Inference-Time Compute via Speculative Reasoning

### v1

@@ Abstract
Recent advances in inference-time compute have significantly improved performance on complex tasks by generating long chains of thought (CoTs) using Large Reasoning Models (LRMs). However, this improved accuracy comes at the cost of high inference latency due to the length of generated reasoning sequences and the autoregressive nature of decoding. Our key insight in tackling these overheads is that LRM inference, and the reasoning that it embeds, is highly tolerant of approximations: complex tasks are typically broken down into simpler steps, each of which brings utility based on the semantic insight it provides for downstream steps rather than the exact tokens it generates. Accordingly, we introduce SpecReason, a system that automatically accelerates LRM inference by using a lightweight model to (speculatively) carry out simpler intermediate reasoning steps and reserving the costly base model only to assess (and potentially correct) the speculated outputs. Importantly, SpecReason’s focus on exploiting the semantic flexibility of thinking tokens in preserving final-answer accuracy is complementary to prior speculation techniques, most notably speculative decoding, which demands token-level equivalence at each step. Across a variety of reasoning benchmarks, SpecReason achieves 1.5–2.5× speedup over vanilla LRM inference while improving accuracy by 1.0–9.9%. Compared to speculative decoding without SpecReason, their combination yields an additional 19.4–44.2% latency reduction. We open-source SpecReason at https://github.com/ruipeterpan/specreason.
@@ 1 Introduction
Inference-time compute has unlocked a new axis in scaling AI progress. The recent advancements of Large Reasoning Models (LRMs) like OpenAI o1/o3 Jaech et al. (2024); ope (2025) and DeepSeek R1 Guo et al. (2025) have demonstrated state-of-the-art accuracy on many complex workloads. Although these LRMs share the same backbone model architecture as traditional Large Language Models (LLMs), their behavior is different: LRMs first “think” by outputting (internal) thinking tokens that break tasks down into composable reasoning steps in a long chain of thought (CoT) Wei et al. (2022) before ultimately outputting tokens that summarize the overall thinking process. Despite their promise, the high quality these models achieve comes at the cost of high inference latency, which is primarily driven by the long reasoning sequences LRMs generate. The problem is further exacerbated by the autoregressive nature of LLMs, which causes latency to scale linearly with sequence length. As a result, final output generation can routinely take minutes — if not hours — to answer a single query; such delays far exceed those from typical LLMs and are prohibitively slow for many interactive applications, ultimately hurting users’ quality of experience Fu et al. (2024b).
Our approach to tackling reasoning delays is rooted in two fundamental properties of LRMs: (1) LRMs approach difficult tasks by generating long CoTs that decompose them into many simpler steps. For example, when solving math problems, a few key reasoning steps require complex long-term planning and greatly influence downstream tasks, while most subsequent steps simply execute the plan through calculations or case analysis (Fig. 1); (2) The utility of a reasoning step hinges less on the exact thinking tokens but more on the semantic insight it contributes to the overall reasoning process – insights that advance the CoT even if the precise wording varies or is imperfect (Fig. 2). Moreover, the self-reflection capabilities of LRMs allow them to correct (occasional) missteps from prior reasoning steps. Taken together, these properties make the decoding of thinking tokens – the primary overhead in LRM inference – inherently more approximation tolerant than typical LLM serving in that a large fraction of intermediate reasoning steps are aligned with the capabilities of, and tolerant of potential errors introduced by, lightweight reasoning models, as our results in Fig. 3 demonstrate.
@@
Building on this observation, we propose SpecReason to accelerate LRM inference while preserving accuracy by carefully offloading these selected, easier intermediate reasoning steps to be speculated by a smaller model. SpecReason uses a smaller reasoning model to produce individual reasoning steps, and only uses the slower but more capable base model to efficiently assess the outputs of the small model (§4.1) and guide reasoning onto the appropriate trajectory (Fig. 1). We find, as prior works have with other tasks Song et al. (2025), that base models can be prompted to act as critic models and judge the utility of intermediate steps well (Fig. 6).
Speculative reasoning vs. speculative decoding. While SpecReason is conceptually similar to speculative decoding Leviathan et al. (2023), an approach that relies on a small draft model to speculate future tokens, there exist key differences. Most notably, speculative decoding is an exact optimization that targets token-level equivalence between the small model and the base model, i.e., focusing on typical LLM serving where all generated tokens are part of the final model output being assessed. In contrast, SpecReason relaxes this stringent requirement and specifically aims to leverage the approximation tolerance that reasoning affords for thinking tokens, i.e., outputs for intermediate reasoning steps. Indeed, we find that semantic-level similarity for thinking tokens (Fig. 2) is enough to preserve final-output accuracy (Fig. 3), which opens up the door for great latency savings in LRM inference. More importantly, because speculative reasoning and speculative decoding operate at different levels, we show that they are complementary optimizations4.2) and can be combined in a hierarchical speculation scheme to achieve even larger latency reductions.
We evaluate SpecReason on a wide range of reasoning workloads that cover common tasks of varying complexity aim (2025); Hendrycks et al. (2021); Rein et al. (2024). Overall, we find that SpecReason reduces end-to-end latency by 1.52.5×1.5-2.5$\times$1.5 - 2.5 × compared to vanilla LRM inference while improving accuracy by $1.0-9.9\%$. Moreover, it is can be combined with speculative decoding to provide a $19.4-44.2\%$ improvement over speculative decoding.
@@ 2 Background
Inference-time scaling. LRMs introduce a structured problem-solving approach that breaks down complex problems into multiple simpler reasoning steps, commonly referred to as a long chain of thought (CoT) Wei et al. (2022). This enables the model to generate intermediate reasoning steps before progressing further and reflect and backtrack to correct errors if needed. LRMs that output long CoTs have been a popular approach to scale inference-time compute Guo et al. (2025); Jaech et al. (2024); ope (2025), and there also exist other schemes like Tree of Thoughts Yao et al. (2023), process-reward-model-guided tree search Lightman et al. (2023); Qi et al. (2024); Guan et al. (2025), and repeated sampling Brown et al. (2024) for scaling inference-time compute.
Speculative decoding. Speculation has long been a classic concept in the literature of computer architecture (Burton, 1985). Due to the memory-bound nature of LLM decoding, recent work has also leveraged the technique of speculation to accelerate the decoding phase (Stern et al., 2018; Leviathan et al., 2023) of LLM inference. The speculative decoding process alternates between speculation and verification steps to ensure correctness while achieving speed-ups. The speculation phase usually consists of either a standalone draft model (Leviathan et al., 2023; Miao et al., 2024), a trainable module on top of the base model (Cai et al., 2024; Li et al., 2025), an n-gram lookup table (Fu et al., 2024a), or a retrieval-based data store (He et al., 2023) to make efficient but less accurate speculations. The verification process, on the other hand, is a base model chunked-prefill over the speculation results, which usually consists of either a single sequence of tokens as in Leviathan et al. (2023) or tree-like structures to further boost the accuracy of speculation (Miao et al., 2024; Cai et al., 2024; Li et al., 2025; Chen et al., 2024). The verification process then accepts the longest matched sequences on the token level from the speculation results and repeats the process. As a result, the speculation length is usually conservative to maintain an optimal trade-off between the speculation overhead and accuracy.
@@ 3 Motivation
Figure 2: The spectrum of approximations of one example reasoning step (equation 1 in Fig. 1).
@@
In this work, we argue that reasoning workloads run on LRMs exhibit unique opportunities for latency reduction thanks to their tolerance to approximations, which sets them apart from traditional generation tasks on LLMs. We illustrate these properties using a representative example from the AIME dataset that was selected for ease of presentation.
@@
Intermediate steps are easier than end-to-end reasoning. One key observation is that the difficulty of reasoning tasks is not uniformly hard across the many reasoning steps within a long CoT. As seen in Fig. 1, while the overall task might be challenging for a small model, only a few key reasoning steps (e.g., analyzing the problem, breaking it down via formulations or case analyses, and coming up with a plan to reason about the problem step-by-step) are critical to the overall reasoning progress, whereas many other reasoning steps are comparatively easy. This behavior is by design: LRMs are trained with RL to generate CoTs that decompose hard problems into a sequence of simpler, more tractable reasoning steps. These intermediate steps often include straightforward reasoning, e.g., arithmetic calculations, case analyses, or logical deductions that are much easier to decode compared to synthesizing the full solution directly in LLM inference. This heterogeneity in difficulty and importance allows a lightweight model to efficiently perform these tasks with sufficient accuracy.
@@
Reasoning progress depends on insights, not exact tokens. Another key takeaway in our work is that the utility of a reasoning step lies not in the exact tokens, but more in the insight it conveys that advances overall reasoning progress. Unlike tasks like translation in traditional LLM inference, where fidelity to exact combinations of tokens matters more, reasoning CoTs within LRM’s thinking tokens care more about the information that advances the reasoning chain. As shown in Fig. 2, there exists a spectrum of valid wordings for a given step: Many semantically-equivalent or semantically-similar sequences of tokens can serve the same functionality in guiding subsequent reasoning.
@@
Occasional mistakes can be corrected via self-reflection. Furthermore, LRMs exhibit strong self-reflection capabilities. Even if an earlier step contains a factual or logical error, LRMs often correct themselves in subsequent reasoning steps, oftentimes marked by outputting tokens like “Wait” or “Hmm”. Moreover, unlike LLM inference where all output tokens are evaluated for accuracy, only the tokens after the thinking tokens count toward accuracy in LRM inference. Therefore, LRM inference can tolerate occasional mistakes, as the models can often identify these mistakes and steer reasoning back onto the correct path during self-reflection. This fault-tolerant property further suggests the potential benefits of approximations.
@@
In summary, compared to traditional LLM inference, LRM inference is inherently more tolerant of approximations that don’t require token-level equivalence as long as the overall reasoning trajectory is maintained. This property is not limited to a single, linear CoT; rather, it extends to more general forms of inference-time compute scaling paradigms, e.g., tree-based searches.
@@ 4.1 Speculative Reasoning
The approximation-tolerant nature of LRM reasoning enables a new form of speculative execution: tentatively carrying out reasoning steps using a lightweight model, assessing their utility with a stronger base model, and selectively accepting them. SpecReason leverages this flexibility to reduce decoding latency while preserving output quality. SpecReason reduces decoding latency by offloading easier or less critical steps – defined as semantically self-contained units like full reasoning steps or sentences – to a smaller, faster model. In decoding each reasoning step, SpecReason proceeds in two stages: (1) the lightweight speculator proposes the next step based on the prior trajectory, and (2) the base model assesses its utility. If accepted, SpecReason moves on to the next step; otherwise, SpecReason falls back to the base model to regenerate the step. While our implementation uses a simple static-threshold verification scheme, the framework supports richer customizations of strategies. We outline key design principles below.
@@
Efficient verification. Because each step requires the base model’s verification, it’s crucial for the verification overhead to remain low to avoid compounding latency. Instead of autoregressively decoding or reranking multiple candidate steps, SpecReason evaluates each speculated step in a single prefill-only pass of the base model. One simple implementation is prompting the base model to generate a single token “score” (e.g., 0-9) that assesses the utility of the speculated step. The assessment prompt is templated to reuse most of the CoT prefix, so each verification requires prefilling only $\sim$70 new tokens. Since short-prefill forward passes are memory-bound, the overhead is comparable to decoding just 1–2 tokens. A static acceptance threshold (e.g., score $\geq 7$) determines whether to accept the speculative step. This threshold can be adjusted to control the aggressiveness of speculation (Fig. 4). If the score falls below the threshold, we discard the speculative step and regenerate it using the base model. We emphasize that this rudimentary static-threshold scheme represents a lower bound on verification quality: more sophisticated alternatives – such as logprobs-based confidence estimates or dynamic thresholds – can be incorporated into SpecReason with no additional cost, and may further improve performance. The key requirement is that verification remains lightweight and step-level.
@@
Implementation details. Due to the lightweight nature of the small model, we colocate the small model and the base model on the same GPU. The memory reserved for KV caches is statically partitioned between the two models. They do not share any internal model states – only the token IDs of the generated reasoning steps are managed and shared by SpecReason. When speculative steps from the small model are rejected, we simply discard the corresponding KV cache entries. Inference is performed sequentially: the small model and base model take turns, avoiding compute kernel interference. In future work, we plan to explore pipelining to overlap the small model’s decoding with the base model’s inference. While this may introduce mild resource contention, it could further reduce end-to-end latency.

### v2

@@ Abstract
Recent advances in inference-time compute have significantly improved performance on complex tasks by generating long chains of thought (CoTs) using Large Reasoning Models (LRMs). However, this improved accuracy comes at the cost of high inference latency due to the length of generated reasoning sequences and the autoregressive nature of decoding. Our key insight in tackling these overheads is that LRM inference, and the reasoning that it embeds, is highly tolerant of approximations: complex tasks are typically broken down into simpler steps, each of which brings utility based on the semantic insight it provides for downstream steps rather than the exact tokens it generates. Accordingly, we introduce SpecReason, a system that automatically accelerates LRM inference by using a lightweight model to (speculatively) carry out simpler intermediate reasoning steps and reserving the costly base model only to assess (and potentially correct) the speculated outputs. Importantly, SpecReason’s focus on exploiting the semantic flexibility of thinking tokens in preserving final-answer accuracy is complementary to prior speculation techniques, most notably speculative decoding, which demands token-level equivalence at each step. Across a variety of reasoning benchmarks, SpecReason achieves 1.43.0×1.4-3.0$\times$1.4 - 3.0 × speedup over vanilla LRM inference while improving accuracy by $0.4-9.0\%$. Compared to speculative decoding without SpecReason, their combination yields an additional $8.8-58.0\%$ latency reduction. We open-source SpecReason at https://github.com/ruipeterpan/specreason.
@@ 1 Introduction
Inference-time compute has unlocked a new axis for scaling AI capabilities. Recent advancements in Large Reasoning Models (LRMs) such as OpenAI o1/o3 (Jaech et al., 2024; ope, 2025) and DeepSeek R1 (Guo et al., 2025) have demonstrated state-of-the-art performance across a wide range of complex tasks. Although these LRMs share the architectural backbones as traditional large language models (LLMs), their inference behavior differs significantly: LRMs first “think” by generating internal thinking tokens—tokens that decompose a task into a sequence of composable reasoning steps via a long chain-of-thought (CoT) (Wei et al., 2022) before producing the final tokens that summarize the reasoning process.
Despite their promise, LRMs incur substantial inference latency due to the length of the reasoning sequences they generate. This challenge is primarily driven by the autoregressive nature of LLMs, where decoding time scales linearly with sequence length. As a result, final output generation can routinely take minutes, if not hours, to answer a single query; such delays far exceed those of typical LLMs and are prohibitively slow for many interactive applications, ultimately degrading user experience (Fu et al., 2024b).
Our approach to tackling reasoning delays—without compromising accuracy—is rooted in two fundamental properties of LRMs: (1) LRMs tackle difficult tasks by generating long CoTs that decompose them into many simpler, sequential steps. For example, in mathematical problem solving, a few key reasoning steps require complex long-term planning and have a major influence on downstream reasoning, while most subsequent steps simply execute the plan through straightforward calculations or case analyses (Fig. 1); (2) The utility of an individual reasoning step hinges less on the exact wording of the thinking tokens but more on the semantic insight it provides. That is, as long as a step contributes meaningfully to advancing the CoT, it remains effective—even if phrased imprecisely or differently (Fig. 2). Moreover, LRMs possess self-reflection capabilities that enable them to revise or correct occasional missteps from earlier steps.
Taken together, these properties make the decoding of thinking tokens—the dominant source of inference latency in LRMs—inherently more approximation tolerant than typical LLM decoding. A large fraction of intermediate reasoning steps can be effectively handled by lightweight reasoning models, which both align with the nature of these steps and can tolerate minor inaccuracies. As shown in Fig. 3, this opens the door to significantly faster inference without sacrificing output quality.
@@
Building on these insights, we propose SpecReason, a system for accelerating LRM inference by selectively offloading easier intermediate steps to be speculated by a smaller model without compromising final output accuracy. SpecReason employs a lightweight reasoning model to generate individual reasoning steps, while reserving the slower but more capable base model to efficiently verify these speculated steps (§4.1) and guide the reasoning process along the correct trajectory (Fig. 1). Consistent with prior findings (Song et al., 2025), we observe that base models can be prompted to act as critic models—assessing the utility of intermediate steps and accepting or rejecting them as needed (Fig. 7).
Speculative reasoning vs. speculative decoding. While SpecReason is conceptually related to speculative decoding (Leviathan et al., 2023), which accelerates LLM inference by using a smaller draft model to predict future tokens, there are key distinctions between the two. Most notably, speculative decoding is an exact optimization: it relies on token-level equivalence between the small and base models, i.e., focusing on typical LLM serving where all generated tokens are part of the final model output being assessed. In contrast, SpecReason explicitly leverages the approximation tolerance inherent in reasoning: it targets thinking tokens—intermediate steps in the reasoning process—where semantic alignment, rather than token-level equivalence, is sufficient. This relaxation enables substantial latency savings during LRM inference, as semantically similar intermediate steps (Fig. 2) are often adequate to preserve end-task accuracy (Fig. 3). In many cases, SpecReason even improves final accuracy over the base model by generating fewer unnecessary tokens (Fig. 4). To further address the high inference cost of LRMs, SpecReason also exposes a user-configurable knob that allows trading off accuracy for latency by adjusting the tolerance level for speculative approximations. Finally and most importantly, because speculative reasoning and speculative decoding operate at different levels, we show that they are complementary techniques (§4.2), and when combined in a hierarchical speculation framework, achieve even greater reductions in inference latency.
We evaluate SpecReason across a wide range of reasoning workloads spanning tasks of varying complexity (aim, 2025; Hendrycks et al., 2021; Rein et al., 2024). Overall, SpecReason reduces end-to-end inference latency by 1.43.0×1.4-3.0$\times$1.4 - 3.0 × compared to vanilla LRM inference while improving accuracy by $0.4-9.0\%$. Moreover, SpecReason can be combined with speculative decoding to provide an additional $8.8-58.0\%$ improvement over speculative decoding alone.
@@ 2 Background
Inference-time scaling. LRMs introduce a structured problem-solving approach that breaks down complex problems into multiple simpler reasoning steps, commonly referred to as a long chain of thought (CoT) (Wei et al., 2022). This enables the model to generate intermediate reasoning steps before progressing further, reflect, and backtrack to correct errors if needed. LRMs that output long CoTs have been a popular approach to scale inference-time compute (Guo et al., 2025; Jaech et al., 2024; ope, 2025), and there also exist other schemes like Tree of Thoughts (Yao et al., 2023), process-reward-model-guided tree search (Lightman et al., 2023; Qi et al., 2024; Guan et al., 2025), and repeated sampling for scaling inference-time compute (Brown et al., 2024).
Speculative decoding. Speculation has long been a classic concept in the literature of computer architecture (Burton, 1985). Due to the memory-bound nature of LLM decoding, recent work has also leveraged the technique of speculation to accelerate the decoding phase (Stern et al., 2018; Leviathan et al., 2023; Yan et al., 2024) of LLM inference. The speculative decoding process alternates between speculation and verification steps to ensure correctness while achieving speed-ups. The speculation phase usually consists of either a standalone draft model (Leviathan et al., 2023; Miao et al., 2024), a trainable module on top of the base model (Cai et al., 2024; Li et al., 2025), a tree-based token cache (Oliaro et al., 2024; Luo et al., 2024; Zhao et al., 2024), an n-gram lookup table (Fu et al., 2024a), or a retrieval-based data store (He et al., 2023) to make efficient but less accurate speculations. The verification process, on the other hand, is a base model chunked-prefill over the speculation results, which usually consists of either a single sequence of tokens as in Leviathan et al. (2023) or tree-like structures to further boost the accuracy of speculation (Miao et al., 2024; Cai et al., 2024; Li et al., 2025; Chen et al., 2024). The verification process then accepts the longest matched sequences on the token level from the speculation results and repeats the process. As a result, the speculation length is usually conservative to maintain an optimal trade-off between the speculation overhead and accuracy.
Existing approaches for reducing latency. Sky-T1-Flash Team (2025) reduces unnecessary thinking tokens by fine-tuning models to curb overthinking, thereby reducing the length of reasoning chains and, consequently, latency. Dynasor-CoT Fu et al. (2024b, 2025) takes a different approach by probing intermediate model confidence and terminating the reasoning process early when the model exhibits sufficient confidence in its current output.
@@ 3 Motivation
Figure 2: The spectrum of approximations of one example reasoning step (equation 1 in Fig. 1). SpecReason can control the exactness of reasoning approximations by adjusting its acceptance threshold to navigate through the accuracy-latency tradeoff space (§5.3).
@@
In this work, we show that reasoning workloads executed by LRMs exhibit unique opportunities for latency reduction due to their inherent tolerance to approximation— setting them apart from traditional generation tasks in LLMs. We illustrate these properties using a representative example from the AIME dataset, selected for its clarity and ease of exposition.
Intermediate steps are easier than end-to-end reasoning. A key observation in LRM behavior is that reasoning difficulty is not uniform across the steps in a long chain-of-thought (CoT). As shown in Fig. 1, while the overall task might be too challenging for a small model to solve end-to-end, only a few critical steps—such as problem analysis, decomposition through formulations or case analyses, and high-level planning—are critical to the overall reasoning progress. In contrast, many other steps are significantly easier.
@@
This behavior is intentional by design: LRMs are often trained with reinforcement learning to generate CoTs that decompose complex problems into sequences of simpler, more tractable reasoning steps. These intermediate steps often include routine reasoning such as arithmetic calculations, case enumeration, or basic logical deductions—operators that are much easier to decode than synthesizing a full solution directly. This heterogeneity in step difficulty and importance creates an opportunity for lightweight models to handle a substantial portion of the reasoning process both efficiently and accurately.
@@
Reasoning progress depends on insights, not exact tokens. Another key takeaway from our work is that the utility of a reasoning step lies in the semantic contribution it makes to the overall reasoning process, rather than the precise tokens it uses. Unlike tasks like translation in traditional LLM inference, where fidelity to exact combinations of tokens matters more, reasoning CoTs within LRM’s thinking tokens care more about the information that advances the reasoning chain. As illustrated in Fig. 2, a spectrum of valid phrasings often exists for a given step: semantically equivalent or similar expressions can convey the same insight and lead to the same downstream reasoning trajectory. This semantic flexibility is a key enabler for approximation-tolerant inference.
@@
Occasional mistakes can be corrected via self-reflection. LRMs exhibit strong self-reflection capabilities, enabling them to recover from earlier reasoning errors. Even when an earlier step contains a factual or logical mistake, the model often revises its trajectory in subsequent steps, marked by tokens like “Wait” or “Hmm”. Moreover, unlike LLM inference where all output tokens contribute to the final answer, in LRM inference, only the tokens generated after the thinking tokens determine the final outcome. Therefore, LRM inference can tolerate occasional mistakes during the reasoning phase, as the model can often identify and correct these mistakes during self-reflection. This inherent fault tolerance further underscores the viability and effectiveness of approximation-based acceleration.
@@
In summary, compared to traditional LLM inference, LRM inference is inherently more tolerant of approximations that do not require token-level equivalence as long as the overall reasoning trajectory is preserved. This property is not limited to a single, linear CoT; rather, it extends naturally to more general inference-time compute scaling paradigms such as tree-based search strategies and other structured reasoning approaches.
@@ 4.1 Speculative Reasoning
Due to its reliance on autoregressive decoding, LRM inference incurs significantly higher latency than typical LLMs—often to the point of being prohibitively slow for interactive applications and degrading user experience (Fu et al., 2025). Existing approaches for latency reduction include using a distilled version of the base model (Guo et al., 2025), limiting the number of thinking tokens via a predefined token budget, or disabling the reasoning process altogether by omitting the thinking tokens (<think> and </think>) during generation (qwe, 2025). However, these approaches impose a harshly trade-off between accuracy for latency: they either limit the model’s capacity to reason or apply a lower-quality model uniformly across all reasoning steps. In contrast, SpecReason takes a more fine-grained and adaptive approach. Instead of explicitly restricting output length, it selectively offloads only the easier reasoning steps to a lightweight model, preserving overall reasoning quality while substantially reducing inference latency.
@@
The approximation-tolerant nature of LRM reasoning enables a new form of speculative execution: tentatively carrying out reasoning steps using a lightweight model, assessing their utility with a stronger base model, and selectively accepting them. SpecReason leverages this flexibility to reduce decoding latency while preserving output quality. To achieve this goal, SpecReason offloads easier or less critical reasoning steps—defined as semantically self-contained units such as complete sentences or logical steps—to a smaller, faster speculator model. Each step is decoded in two stages: (1) the lightweight speculator proposes the next reasoning step based on the current context, and (2) the base model evaluates the proposed step for semantic utility. If the step is accepted, SpecReason proceeds to the next step; otherwise, SpecReason falls back to the base model to regenerate the step. While our implementation uses a simple static-threshold mechanism for verification, the framework supports richer, customizable decision strategies. We outline key design principles below.
Navigating the Pareto frontier of the latency-accuracy tradeoff. SpecReason expands the Pareto frontier of the latency-accuracy tradeoff by exposing fine-grained control knobs to navigate through this space. The key knob SpecReason employs is the acceptance threshold: after each speculated reasoning step, the base model is prompted to generate a single-token utility score (e.g., an integer from 0 to 9) indicating the quality of the step. If the utility score is above a static acceptance threshold (e.g., score $\geq 7$), the speculated reasoning step is accepted; otherwise, it is discarded and regenerated by the base model.
Adjusting this threshold allows users to control the strictness of speculation (Fig. 5): a higher threshold requires speculated steps to be closer to token-level equivalence on the equivalence spectrum (Fig. 2), improving accuracy but reducing the acceptance rate and thereby increasing latency. Conversely, a lower threshold increases speculation efficiency at the cost of potential accuracy degradation.
An additional knob involves forcing the first $n$ reasoning steps to be decoded by the base model. Since LRMs often use the initial steps to analyze the problem and formulate a high-level plan, assigning these initial steps to the base model can steer the overall reasoning trajectory toward higher quality. We show in Fig. 7 that this knob also allows SpecReason to manage latency-accuracy tradeoff, though with less impact than the acceptance threshold knob.
While our current implementation uses a simple, discrete threshold-based scoring scheme—offering only a coarse-grained configuration space—it establishes a lower bound on verification quality. Future work can explore more sophisticated strategies, such as logprob-based confidence estimates or dynamic thresholds, to enable finer-grained tradeoffs without incurring additional runtime cost, and may further improve overall performance.
@@
Efficient verification. Because each step requires verification by the base model, it’s crucial to keep verification overhead low to avoid compounding latency. Instead of autoregressively decoding or reranking multiple candidate steps, SpecReason evaluates each speculated step in a single prefill-only pass of the base model. The verification prompt is templated to reuse most of the CoT prefix, so each verification requires prefilling only $\sim$70 new tokens. Since short-prefill forward passes are memory-bound, the overhead is comparable to decoding just 1–2 tokens, making verification highly efficient in practice.
@@
Implementation details. Since the small model is lightweight, we colocate both the small and base models on the same GPU. The memory reserved for Key-Value caches (Kwon et al., 2023) is statically partitioned between the two models. They do not share any internal model states–only the token IDs of the generated reasoning steps are managed and shared by SpecReason. If a speculative step is rejected, the corresponding KV cache entries are discarded.
Inference is performed sequentially: the small and base models take turns, avoiding kernel-level interference. In future work, we plan to explore pipelining to overlap the small model’s decoding with the base model’s inference. While this may introduce mild resource contention, it could further reduce end-to-end latency.